Find four consecutive even integers so that the sum of the first three is 2 more than twice the fourth.: Find four consecutive even integers so that the sum of the first three is 2 more than twice the fourth.
Answer by stanbon(18998)

Find four consecutive even integers so that the sum of the first three is 2 more than twice the fourth.
------------------------------------
1st: 2x
2nd: 2x + 2
3rd: 2x + 4
4th: 2x + 6
-----------------
EQUATION:
2x + 2x+2 + 2x+6 = 2(2x+6) + 2
6x + 8 = 4x + 14
2x = 6
x = 3
-------------
1st: 2x = 6
2nd: 2x+2 = 8
3rd: 2x+4 = 10
4th: 2x+6 = 12
==============
Cheers,
Stan H.